OpenStudy (anonymous):

Solve: |3r - 12| = 0 A. {-4, 4} B. {0, 4} C. {0, -4} D. {4}

OpenStudy (callisto):

|3r - 12| = 0 3r - 12= 0 -(3r - 12) = 0 3r =12 or -3r =-12 Can you solve it here?